3x-square root of (3x+7) = -5
Answer by stanbon(75887)

3x-square root of (3x+7) = -5
------
3x - sqrt(3x+7) = -5
-----
sqrt(3x+7) = 3x+5
-----
3x+7 = 9x^2+30x+25
-----
9x^2 + 27x + 18 = 0
----
x^2 + 3x + 2 = 0
Factor:
(x+2)(x+1) = 0
--------
x = -1 or x = -2
===================
Cheers,
Stan H.
===================